# DepthCrafter: Generating Consistent Long Depth Sequences for Open-world Videos

### v1

@@ 3.3 Training
DepthCrafter (Ours) 0.292 0.697 0.125 0.848 0.110 0.881 0.075 0.971 0.076 0.944
Table 1: Zero-shot relative depth estimation results. We compare with the best single-image depth estimation model, Marigold [32], and Depth Anything (V2) [67, 68], as well as the representative video depth estimation models, NVDS [63] and ChronoDepth [53]. Best and second best results are highlighted.
3.4 Inference for Extremely Long Videos
4 Experiments
4.1 Implementation
We implemented our DepthCrafter based on SVD [3], using the diffusers [58] library. We train our model at the resolution of $320\times 640$ for efficiency, but we can estimate depth sequences at any resolution, e.g., $576\times 1024$, during inference. We use the Adam optimizer [34] with a learning rate of $1\times 10^{-5}$ and a batch size of 8. The number of iterations in the three stages of training is $80K$, $40K$, and $10K$, respectively. We employed eight NVIDIA A100 GPUs for training, with a total training time of about five days. We also adopt the classifier-free guidance [20] to improve the details of the generated depth sequences. The number of denoising steps is set to 25 for all experiments.
4.2 Evaluation
Evaluation datasets. We evaluate our model on four video datasets, a single-image dataset, as well as the DAVIS dataset [45] and in-the-wild videos for qualitative results. All the evaluation videos were not included in our training process. Sintel [6] is a synthetic dataset with precise depth labels, featuring dynamic scenes with diverse content and camera motion. It contains 23 sequences with the length of around 50 frames each in the training set. ScanNet v2 [11] is an indoor dataset with depth maps obtained from a Kinect sensor. For evaluation purposes, we employed the test set, which includes 100 RGB-D video sequences of various scenes. We extracted 90 frames from each sequence at a rate of 15 frames per second. Since ScanNet v2 contains only static indoor scenes, we further introduced 5 dynamic indoor RGB-D videos with a length of 110 frames each from the Bonn [43] dataset to better evaluate the performance of our model on dynamic scenes. KITTI [18] is a street-scene outdoor dataset for autonomous driving, with sparse metric depths captured by a LiDAR sensor. We adopted the validation set, which includes 13 scenes, and extracted 13 videos from it with a length of 110 frames each. Besides, we also evaluated our model for single-image depth estimation on the NYU-v2 [54] dataset, which contains 654 images in the test split. These datasets cover a wide range of scenes, including synthetic and realistic scenes, indoor and outdoor scenes, and static and dynamic scenes, to evaluate the generalization ability of our model across various open-world scenarios.
Figure 4: Qualitative comparison for open-world video depth estimation. We compare with the representative single-image depth estimation method Depth-Anything-V2 [68] and the video depth estimation method NVDS [63]. For better visualizing the temporal quality, we show the temporal profiles of each result in red boxes, by slicing the depth values along the time axis at the red line positions.
@@ 4.2 Evaluation
Quantitative results. We compare our DepthCrafter with the representative methods for both single-image and video depth estimation, i.e. Marigold [32], Depth-Anything [67], Depth-Anything-V2 [68], NVDS [63], and ChronoDepth [53]. As shown in Tab. 1, our DepthCrafter achieves state-of-the-art performance in all four video datasets, thanks to the powerful open-world video undederstanding capability of the video diffusion models and the three-stage training strategy that leverages both realistic and synthetic datasets. For Sintel and KITTI, characterized by significant camera motion and fast-moving objects, our DepthCrafter outperforms the previous strongest Depth-Anything (V2) model tremendously in terms of both the AbsRel and $\delta_{1}$ metrics, e.g. $(0.697-0.564)/0.564=23.6\%$ improvement in $\delta_{1}$ on Sintel. For indoor datasets like Scannet and Bonn, featuring minimal camera motion and roughly the same room scales, Depth-Anything has exhibited strong performance. Nevertheless, we still have some performance enhancements over Depth-Anything, e.g. $(0.130-0.125)/0.130=3.8\%$ improvement in AbsRel on Scannet. Note that the sequence length of these datasets varies from 50 to 110 frames, and our model can generalize well across different video lengths.
Qualitative results. To further demonstrate the effectiveness of our model, we present the qualitative results on video depth estimation from the DAVIS dataset [45], Sora generated videos [5], and open-world videos, including human actions, animals, architectures, cartoons, and games, where the sequence length varies from 90 to 195 frames. As shown in Fig. 4, we show the temporal profiles of the estimated depth sequences in the red line position by slicing the depth values along the time axis, to better visualize the temporal consistency of the estimated depth sequences, following the practice in [63, 25]. We can observe that our DepthCrafter can produce temporally consistent depth sequences with fine-grained details across various open-world videos, while both NVDS and Depth-Anything exhibit zigzag artifacts in the temporal profiles, indicating the flickering artifacts in the estimated depth sequences. These results demonstrate the effectiveness of our DepthCrafter in generating temporally consistent long depth sequences with high-fidelity details for open-world videos.
@@
Figure 5: Results on depth estimation from static images. We compare with the representative single-image depth estimation method Depth-Anything-V2 [68].
@@ 4.3 Ablation Studies
Effectiveness of the three-stage training strategy. We first ablate the effectiveness of the three-stage training strategy by evaluating the performance of our model at the end of each stage on the Sintel dataset [6], since it contains precise depth annotations on dynamic scenes. From Tab. 2, we can observe that the performance of our model almost improves as the training progresses, indicating the effectiveness of the three-stage training strategy. Although the AbsRel metric slightly increases in stage 2, the $\delta_{1}$ metric consistently improves, and stage 2 is essential for supporting the long temporal context up to 110 frames.
Effectiveness of the inference strategy. To ablate the effectiveness of our inference strategy components, we consider these variants: baseline, which independently infers each segment and directly averages the overlapped frames; + initialization, which contains the same initialization of overlapped latents as our method, but without the stitching process; + initialization & stitching, which is our full method. We visually compare the temporal profiles of the estimated depth sequences of these variants in Fig. 6. We can observe the overlapped jaggies in both the static regions (pointed by the yellow arrow) and the dynamic regions (pointed by the green arrow) in temporal profiles of the “baseline” method, which indicates the flickering artifacts. The “+ initialization” method can alleviate the flickering artifacts in the static regions, but still has jaggies in the dynamic regions, while our full method can produce smooth depth sequences in both static and dynamic regions.
stage 1 stage 2 stage 3
AbsRel $\downarrow$ 0.323 0.336 0.292
$\delta_{1}$uparrow 0.625 0.673 0.697
@@ 4.4 Applications
Our DepthCrafter can facilitate various downstream applications, e.g., foreground matting, depth slicing, fog effects, and depth-conditioned video generation, by providing temporally consistent depth sequences with fine-grained details for open-world videos. We show example results of fog effects and depth-conditioned video generation in Fig. 7, while more visual effects results are available in our website. For the fog effect, we blend the fog map with the input video frames based on the depth values to simulate varying transparency levels in fog. And many recent conditioned video generation models [9, 19, 15, 75] employ depth maps as the structure conditions for video generation or editing. We adopt Control-A-Video [9] and video depth of our method as conditions to generate a video with prompts “a rider walking through stars, artstation”. The visual effects of these applications rely heavily on the accuracy and consistency of the video depth, which demonstrates the wide applicability of our DepthCrafter in various downstream tasks.
Figure 6: Ablation studies on the effectiveness of the inference strategy. We profile the estimated depth sequences of different variants on the red line position. The yellow and green arrows point to the static and dynamic regions, respectively.
@@ 5 Conclusion
We present DepthCrafter, a novel method for open-world video depth estimation by leveraging video diffusion models. It can generate temporally consistent depth sequences with fine-grained details for video width diverse content, motion, and camera movement, without requiring any additional information. It also supports videos of variable lengths, ranging from one frame (static image) to extremely long videos. This is achieved through our meticulously designed three-stage training strategy, compiled paired video-depth datasets, and an inference strategy. Extensive evaluations have demonstrated that DepthCrafter achieves state-of-the-art performance in open-world video depth estimation under zero-shot settings. It also facilitates various downstream applications, including depth-based visual effects and conditional video generation. There are still some limitations to be addressed in the future, such as the expensive computation and memory cost, which is due to the large model size and the iterative denoising process in the diffusion model.

### v2

@@ 4.2 Evaluation
DepthCrafter (Ours) 0.270 0.697 0.123 0.856 0.104 0.896 0.071 0.972 0.072 0.948
@@
Quantitative results. We compare our DepthCrafter with the representative methods for both single-image and video depth estimation, i.e. Marigold [32], Depth-Anything [67], Depth-Anything-V2 [68], NVDS [63], and ChronoDepth [54]. For Depth-Anything (V2), we use the large model variant for its best performance, while for Marigold we adopt the LCM version with an ensemble size of 5. As shown in Tab. 1, our DepthCrafter achieves state-of-the-art performance in all four video datasets, thanks to the powerful open-world video undederstanding capability of the video diffusion models and the three-stage training strategy that leverages both realistic and synthetic datasets. For Sintel and KITTI, characterized by significant camera motion and fast-moving objects, our DepthCrafter outperforms the strong Depth-Anything (V2) model tremendously in terms of both the AbsRel and $\delta_{1}$ metrics, e.g. $25.7\%=(0.140-0.104)/0.140$ improvement in AbsRel on KITTI. For indoor datasets like Scannet and Bonn, featuring minimal camera motion and roughly the same room scales, Depth-Anything has exhibited good performance. Nevertheless, we still have performance enhancements over it, e.g. $5.4\%=(0.130-0.123)/0.130$ improvement in AbsRel on Scannet. Note that the sequence length of these datasets varies from 50 to 110 frames, and our model can generalize well across different video lengths.
Figure 5: Single-image depth estimation. We compare with the representative single-image depth estimation method Depth-Anything-V2 [68].
Qualitative results. To further demonstrate the effectiveness of our model, we present the qualitative results from the DAVIS dataset [46], Sora generated videos [6], and open-world videos, including human actions, animals, architectures, cartoons, and games, where the sequence length varies from 90 to 195 frames. As shown in Fig. 4, we show the temporal profiles of the estimated depth sequences in the green line position by slicing the depth values along the time axis, to better visualize the temporal consistency of the estimated depth sequences, following the practice in [63, 26]. We can observe that our DepthCrafter can produce temporally consistent depth sequences with fine-grained details across various open-world videos, while both NVDS and Depth-Anything exhibit zigzag artifacts in the temporal profiles, indicating the flickering artifacts in the estimated depth sequences. These results demonstrate the effectiveness of our DepthCrafter in generating temporally consistent long depth sequences with high-fidelity details for open-world videos.
@@
stage 1 stage 2 stage 3
AbsRel $\downarrow$ 0.322 0.316 0.270
$\delta_{1}$uparrow 0.626 0.675 0.697
@@ 4.3 Ablation Studies
Effectiveness of the three-stage training strategy. We first ablate the effectiveness of the three-stage training strategy by evaluating the performance of our model at the end of each stage on the Sintel dataset [7], since it contains precise depth annotations on dynamic scenes. From Tab. 2, we can observe that the performance of our model consistently improves as the number of training stages increases, indicating the effectiveness of the three-stage training strategy. More ablation results on other datasets are available in the supplementary material.
Effectiveness of the inference strategy. To ablate the effectiveness of our inference strategy components, we consider these variants: baseline, which independently infers each segment and directly averages the overlapped frames; + initialization, which contains the same initialization of overlapped latents as our method, but without the stitching process; + initialization & stitching, which is our full method. We visually compare the temporal profiles of the estimated depth sequences of these variants in Fig. 6. We can observe the overlapped jaggies in both the static regions (pointed by the yellow arrow) and the dynamic regions (pointed by the blue arrow) in temporal profiles of the “baseline” method, which indicates the flickering artifacts. The “+ initialization” method can alleviate the flickering artifacts in the static regions, but still has jaggies in the dynamic regions, while our full method can produce smooth depth sequences in both static and dynamic regions.
Figure 6: Ablation studies on the effectiveness of the inference strategy. We profile the estimated depth sequences of different variants on the green line position. The yellow and blue arrows point to the static and dynamic regions, respectively.
@@ 4.4 Applications
Our DepthCrafter can facilitate various downstream applications, e.g., foreground matting, depth slicing, fog effects, and depth-conditioned video generation, by providing temporally consistent depth sequences with fine-grained details for open-world videos. We show example results of fog effects and depth-conditioned video generation in Fig. 7, while more visual effects results are available in the supplementary material. For the fog effect, we blend the fog map with the input video frames based on the depth values to simulate varying transparency levels in fog. And many recent conditioned video generation models [10, 20, 16, 76] employ depth maps as the structure conditions for video generation or editing. We adopt Control-A-Video [10] and video depth of our method as conditions to generate a video with prompts “a rider walking through stars, artstation”. The visual effects of these applications rely heavily on the accuracy and consistency of the video depth, which demonstrates the wide applicability of our DepthCrafter in various downstream tasks.
@@
4.5 Limitations
Although our DepthCrafter achieves state-of-the-art performance in open-world video depth estimation, there are still some limitations to be addressed in the future. First, the computation and memory consumption of our model is relatively high, which is mainly due to the large model size and the iterative denoising process in the diffusion model. As shown in Tab. 3, we report the inference speed of our model, Depth-Anything (V2), and Marigold, with the resolution of $1024\times 576$ on a single NVIDIA A100 GPU. We can see that Depth-Anything (V2) is faster than our model, while our method is much faster than Marigold since it ensembles five inference results. Our model achieves the best quality without ensembling, and its inference speed of 465.84 ms per frame is acceptable for many applications. Second, our method requires around 24GB GPU memory to process a video with the resolution of $1024\times 576$ and a segment length of 110 frames. But note that, with the segment length of 40 frames, the memory consumption can be reduced to 12GB, which is acceptable for most modern GPUs, and we can still infer videos of any length by our inference strategy under this setting. We believe that further engineering efforts, such as model distillation and quantization, can help further promote the practicality of our method.
Table 3: Inference time per frame (ms) of our model, Depth-Anything (V2), and Marigold, with the resolution of $1024\times 576$.
Method Encoding Denoising Decoding All
Depth-Anything (V2) N/A N/A N/A 180.46
Marigold 256.40 114.53 699.36 1070.29
DepthCrafter (Ours) 51.85 160.93 253.06 465.84
@@ 5 Conclusion
We present DepthCrafter, a novel open-world video depth estimation method that leverages video diffusion models. It can generate temporally consistent depth sequences with fine-grained details for video width diverse content, motion, and camera movement, without requiring any additional information. It also supports videos of variable lengths, ranging from one frame (static image) to extremely long videos. This is achieved through our meticulously designed three-stage training strategy, compiled paired video-depth datasets, and an inference strategy. Extensive evaluations have demonstrated that DepthCrafter achieves state-of-the-art performance in open-world video depth estimation under zero-shot settings. It also facilitates various downstream applications, including depth-based visual effects and conditional video generation.
Supplementary Material
Appendix
In this supplementary material, we provide additional implementation details in Appendix A and more evaluations in Appendix B. For more visual results and details, we highly recommend referring to the webpage: https://depthcrafter.github.io, since the visual quality of the generated depth sequences can be better accessed with interactive videos. We would release the code and model to facilitate further research and applications.
Appendix A Implementation Details
A.1 Data Preparation
Following conventional practice in depth estimation [67, 68], we represent the depth in the disparity domain. We target relative depth estimation, so we normalize the disparity values to the range of $[0,1]$ by the maximum and minimum disparity values in the sequence. Since the training of our DepthCrafter only involves the U-Net model, with the VAE freezing, we can pre-process the latents of videos and the corresponding depth sequences in advance. This caching mechanism significantly reduces the training time and memory consumption, as the latents do not need to be re-computed during the training process, and the VAE does not need to be loaded into the memory.
A.2 Training Details
We follow the EDM-framework [31] to train our DepthCrafter. This can be mathematically formulated as Eqs. 1, 2 and 3. The $c_{\text{in}}$, $c_{\text{out}}$, $c_{\text{skip}}$, and $c_{\text{noise}}$ in Eq. 3 are the EDM preconditionining functions [31, 53]:
$\begin{split}c_{\text{in}}(\sigma_{t})&=1/\sqrt{1+\sigma_{t}^{2}},\\ c_{\text{out}}(\sigma_{t})&=-\sigma_{t}/\sqrt{1+\sigma_{t}^{2}},\\ c_{\text{skip}}(\sigma_{t})&=1/(1+\sigma_{t}^{2}),\\ c_{\text{noise}}(\sigma_{t})&=0.25\cdot\log(\sigma_{t}).\end{split}$ (5)
$c_{\text{in}}$ and $c_{\text{out}}$ are used to scale the input and output magnitudes, $c_{\text{skip}}$ is used to modulate the skip connection, and $c_{\text{noise}}$ is used to map the noise level $\sigma_{t}$ into a conditioning input for the denoiser $F_{\theta}$. The $\lambda_{\sigma_{t}}$ in Eq. 2 effectively incurs a per-sample loss weight for balancing different noise levels, which is set as:
$\lambda_{\sigma_{t}}=1/c_{\text{out}}(\sigma_{t})^{2}.$ (6)
During training, we randomly sample the noise level $\sigma_{t}$ from a log-normal distribution:
$\text{ln}(\sigma_{t})\sim\mathcal{N}(0.7,1.6^{2}),$ (7)
which is following the EDM-framework [31] to target the training efforts to the relevant range.
We train our DepthCrafter on eight NVIDIA A100 GPUs with a learning rate of $10^{-5}$, and a batch size of 8. We adopted the DeepSpeed ZeRO-2 strategy, gradient checkpointing, and mixed precision training to reduce memory consumption during training. We also highly optimize the U-Net structure and cache the latents to further reduce memory consumption. The first and third training stages consume around 40GB of GPU memory per device, while the second stage consumes around 80GB. The “temporal layers” mentioned in Sec. 3 are the layers performed on the time axis, such as temporal transformer and temporal resnet blocks. The remaining layers are spatial layers, such as spatial transformers and spatial resnet blocks.
Figure S1: Effects of the number of denoising steps in our DepthCrafter. We show an example from the Bonn dataset [44], where the depth sequences are generated with different numbers of denoising steps. The green and orange arrows indicate the regions where more denoising steps can refine the depth details.
Table S1: Performance comparison of our DepthCrafter with different numbers of denoising steps. For reference, we also include the results of Marigold [32] and Depth-Anything-V2 [68]. The inference speed is measured in milliseconds per frame at the resolution of 1024$\times$576. Best and second best results are highlighted.
Method Steps ms / frame $\downarrow$ Sintel ($\sim$50 frames) Scannet (90 frames) KITTI (110 frames) Bonn (110 frames)
[…]
Marigold [32] 1070.29 0.532 0.515 0.166 0.769 0.149 0.796 0.091 0.931
Depth-Anything-V2 [68] 180.46 0.367 0.554 0.135 0.822 0.140 0.804 0.106 0.921
DepthCrafter (Ours) 1 337.10 0.319 0.651 0.132 0.826 0.138 0.812 0.084 0.954
2 369.28 0.301 0.661 0.132 0.828 0.138 0.814 0.083 0.955
3 401.47 0.273 0.693 0.123 0.854 0.111 0.877 0.073 0.971
4 433.65 0.293 0.697 0.123 0.856 0.107 0.888 0.072 0.971
5 465.84 0.270 0.697 0.123 0.856 0.104 0.896 0.071 0.972
6 498.03 0.299 0.696 0.124 0.851 0.105 0.891 0.072 0.973
10 626.72 0.291 0.694 0.125 0.849 0.106 0.890 0.073 0.972
25 1109.42 0.292 0.697 0.125 0.848 0.110 0.881 0.075 0.971
Table S2: Effectiveness of our three-stage training strategy. We show the performance of our DepthCrafter with different training stages on the Sintel, Scannet, KITTI, and Bonn datasets. For reference, we also include the results of Marigold [32] and Depth-Anything-V2 [68]. Best and second best results are highlighted.
Method Training Stages Sintel ($\sim$50 frames) Scannet (90 frames) KITTI (110 frames) Bonn (110 frames)
[…]
Marigold [32] 0.532 0.515 0.166 0.769 0.149 0.796 0.091 0.931
Depth-Anything-V2 [68] 0.367 0.554 0.135 0.822 0.140 0.804 0.106 0.921
DepthCrafter (Ours) 1 0.322 0.626 0.170 0.721 0.174 0.724 0.103 0.917
2 0.316 0.675 0.134 0.826 0.127 0.844 0.090 0.935
3 0.270 0.697 0.123 0.856 0.104 0.896 0.071 0.972
A.3 Benchmark Evaluation Details
Since existing monocular depth estimation methods and benchmarks are mainly tailored for static images, we re-compile the public benchmarks to evaluate the video depth estimation methods. First, we re-format the testing datasets in the form of videos that are originally in the form of images. Specifically, for the ScanNet V2 dataset [12], we extracted the first 90 RGB-D frames from the original sensor data sequences at a rate of 15 frames per second. Besides, there are black regions in the corners of the images due to the camera calibration, which would affect the depth estimation evaluation. We carefully crop the borders of the images to remove the black regions, i.e. cropping 8 pixels from the top and bottom, and 11 pixels from the left and right. For the KITTI dataset [19], we extracted the first 110 frames from the original sequences without downsampling the frame rate, since the difference between consecutive frames is relatively large. And for the Bonn dataset [44], which was usually not included in the evaluation due to the small scale, but we find it is a good complement to the ScanNet dataset for indoor scenes as it contains dynamic contents while ScanNet contains only static scenes. We selected five sequences from the Bonn dataset, each with 110 frames, for evaluation. For the Sintel dataset [7], as it is a synthetic dataset, we directly used the original sequences.
For the evaluation metrics, we followed the insight from the commonly used metrics in the depth estimation, including the absolute relative error (AbsRel) and the $\delta_{1}$ metric, but modified the scale and shift alignment from per-image to per-video. This is because the depth values for a video should be consistent across frames, otherwise, the depth sequences would be flickering. During evaluation, we first align the depth sequences to the ground truth by the scale and shift, using a least-square optimization. Following MiDas [49], we cap the maximum depth values to a certain value for different datasets, e.g., 70 meters for the SinTel dataset, 80 meters for the KITTI dataset, and 10 meters for the ScanNet, Bonn, and NYUv2 datasets.
Figure S2: Effects of the classifier-free guidance (CFG) on the generated depth sequences. For better visualization, we blow up two regions that contain structures with fine-grained details.
Table S3: Effects of the classifier-free guidance (CFG). For reference, we also include the results of Marigold [32] and Depth-Anything-V2 [68]. Best and second best results are highlighted.
Method Sintel ($\sim$50 frames) Scannet (90 frames) KITTI (110 frames) Bonn (110 frames)
[…]
Marigold [32] 0.532 0.515 0.166 0.769 0.149 0.796 0.091 0.931
Depth-Anything-V2 [68] 0.367 0.554 0.135 0.822 0.140 0.804 0.106 0.921
DepthCrafter W/O CFG 0.270 0.697 0.123 0.856 0.104 0.896 0.071 0.972
DepthCrafter W/ CFG 0.315 0.692 0.123 0.850 0.108 0.885 0.076 0.972
Appendix B Additional Evaluations
B.1 Effect of Number of Denoising Steps
During inference, the number of denoising steps is a crucial hyperparameter that affects the trade-off between the inference speed and the depth estimation quality. The practice in image-to-video diffusion models [3] is to set the number of denoising steps to around 25. However, as shown in Fig. S1, we find that the number of denoising steps can be reduced significantly for video depth estimation, even one step works well. This is because the video depth estimation task is more deterministic than the video generation task. And we can see in the figure that more denoising steps would consistently improve the structure details of the generated depth sequences. In Tab. S1, we show the results of our DepthCrafter with different numbers of denoising steps. We can see that our DepthCrafter significantly outperforms existing strong baselines, such as Marigold [32] and Depth-Anything-V2 [68], even with only one denoising step. The performance of our DepthCrafter is increased with more denoising steps, but the improvement gets saturated after five steps. Thus we set the number of denoising steps to five in our experiments, which achieves a good trade-off between the inference speed and the depth estimation quality. The inference speed of our DepthCrafter with five denoising steps is 465.84 ms per 1024$\times$576 frame, which is acceptable for many applications.
B.2 Effectiveness of Training Stages
In the main paper, we ablate the performance of our DepthCrafter with three training stages, only on the Sintel [7] dataset. To complement the evaluation, we further evaluate the effectiveness of our three-stage training strategy on all the datasets, including Sintel [7], Scannet [12], KITTI [19], and Bonn [44]. As shown in Tab. S2, we can observe that, even only with the first two stages, our DepthCrafter already outperforms the existing strong baselines, such as Marigold [32] and Depth-Anything-V2 [68]. More importantly, the performance improvement with the training stages is consistent across all the datasets. It indicates that our three-stage training strategy is effective for improving the generalization ability of our DepthCrafter to diverse open-world videos.
B.3 Effects of Classifier-Free Guidance
Classifier-free guidance (CFG) is proven to be effective in improving the details of the generated videos in video diffusion models [3, 8, 9, 65, 64]. In our DepthCrafter, we also investigate the effectiveness of CFG in video depth estimation. As shown in Fig. S2, we show an example frame from the KITTI dataset, where the results of our DepthCrafter with and without CFG are compared. We can see that the CFG can indeed improve the visual details of the generated depth sequences, especially for the fine-grained structures. However, we find that the CFG may slightly degrade the quantitative accuracy of the depth estimation, as shown in Tab. S3. This may be because the CFG is designed for improving the details of the generated videos, while the depth estimation task is more deterministic and requires more accurate predictions. Since adopting the CFG would also introduce additional computation, we do not use the CFG in our DepthCrafter for the main experiments. However, if the users are more interested in the visual details of the depth sequences, they can consider incorporating the CFG into our DepthCrafter.